A researcher wants to estimate the proportion of the population which possesses a given characteristic. A random sample of size 1800 is taken resulting in 450 items which possess the characteristic. The point estimate for this population proportion is _______.
Answer #1

Point of eatimest, p = x / n = 450/1800 = 0.25 (ans)